OpenStudy (anonymous):

Find all solutions to the equation. 7 sin^2x - 14 sin x + 2 = -5

OpenStudy (misty1212):

HI!!

OpenStudy (anonymous):

let sin x=t

OpenStudy (misty1212):

solve \[7u^2-14u-7=0\]

OpenStudy (anonymous):

hello

OpenStudy (misty1212):

then replace \(u\) by \(\sin(x)\)

OpenStudy (anonymous):

7t^2-14t+7=0 t^2-2t+1=0

OpenStudy (anonymous):

solve it as a regular quadratic equation

OpenStudy (anonymous):

(t-1)^2=0 t=+1,-1

OpenStudy (anonymous):

case 1: sin x=1 case 2: sin x=-1

OpenStudy (misty1212):

yeah i made a mistake, should be \[7u^2-14u+7=0\] but the above answer is incorrect

OpenStudy (anonymous):

so what would the solutions be?

OpenStudy (anonymous):

oh sorry sin x=-1 cant be possible only sin x=1

OpenStudy (misty1212):

\[(u-1)^2=0\\ u=1\\ \sin(x)=1\] etc

OpenStudy (anonymous):

those all the solutions?

OpenStudy (anonymous):

\[\sin(x)=\pm1\] CASE 1 \[\sin(x)=1\] \[x=(4n+1)\frac{\pi}{2}, n \in \mathbb{Z}\]

OpenStudy (anonymous):

similarly solve for sin(x)=-1

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

You have to solve for x, when sin(x)=1, x can take the values \[\frac{\pi}{2},\frac{5\pi}{2},\frac{9\pi}{2},.....\]\[\frac{\pi}{2},2\pi+\frac{\pi}{2},4\pi+\frac{\pi}{2},.....\]\[(2\times 0 \times \pi)+\frac{\pi}{2},(2 \times 1 \times \pi)+\frac{\pi}{2},(2 \times 2 \times \pi)+\frac{\pi}{2}\]\[2n \pi+\frac{\pi}{2},n \in \mathbb{Z}\]Thus there are infinite solutions, for any integral value of n\[-3,-2,-1,0,1,2,3\]

OpenStudy (anonymous):

which is the same as \[(4n+1)\frac{\pi}{2}=2n \pi+\frac{\pi}{2}\]

OpenStudy (anonymous):

@ganeshie8